OpenStudy (anonymous):

Find the direct variation equation of the graph through the points (0, 0) and (1, -2). Write in y=kx form. y = 2x y = -2x y = 1/2 x y = -1/2 x

OpenStudy (anonymous):

guess i didn't help the first time. do you know what exactly you are looking for in this problem?

OpenStudy (anonymous):

you are looking for a number, that they call "\(k\)" in this problem your choices from the possible answers are either \(k=2,k=-2,k=\frac{1}{2}\) or \(k=-\frac{1}{2}\) and your job it to find the correct one

OpenStudy (anonymous):

i just dont know how to change the form

OpenStudy (anonymous):

the point \((0,0)\) is not much help to solve this, but the point \((1,-2)\) is because this tells you that if \(x=1\) then \(y=-2\) is that clear or not?

OpenStudy (anonymous):

in other words, you can find \(k\) by looking at the equation \[y=kx\] replace \(x\) by \(1\) and \(y\) by \(-2\)

OpenStudy (anonymous):

the first numkber or x is 1 and the second number or y is -2?

OpenStudy (anonymous):

so -2=k1?

OpenStudy (anonymous):

yes in that particular example \(x\) and \(y\) are variables, but you know that IF \(x=1\) THEN \(y=-2\)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

and since \(k\times 1=k\) you know that \(-2=k\) which is what you were looking for

OpenStudy (anonymous):

so now go back the the original equation \[y=kx\] and where you see a "\(k\)" replace it by \(-2\)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

it im a little confused how do you end uo with k= instead of y=

OpenStudy (anonymous):

you would write \[1=k\times 4\] and solve for \(k\) by dividing both sides by \(4\) to get \[\frac{1}{4}=k\] then go back to \[y=kx\] and replace \(k\) by \(\frac{1}{4}\)

OpenStudy (anonymous):

yes i see that you are confused by that \(k\) is some fixed number that you are trying to find \(x\) and \(y\) are variables

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

so for example, once we know the equation is \[y=-2x\] for the first example, \(x\) and \(y\) can vary, they are not fixed numbers for example if \(x=3\) then \(y=-2\times 3=-6\) and if \[x=5\] then \[y=-2\times 5=-10\]

OpenStudy (anonymous):

the points would look like \((3,-6)\) and \((5,-10)\)